OpenStudy (anonymous):

Identify the point (x1, y1) from the equation: y + 8 = 3(x – 2) (–2, 8) (8, –2) (–8,2) (2, –8)

OpenStudy (campbell_st):

you have the point slope form of a straight line equation so in general its \[y - y_{1} = m (x - x_{1})\] where \[m = slope...and.... (x_{1}, y_{1}) = point\] match this to your equation and you should be able to identify the point

OpenStudy (kkutie7):

The last one